4. Which term of the A.P. 3, 8, 13, 18, … is 78?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given the A.P. series as3, 8, 13, 18, … First term, a = 3 Common difference, d = a2 − a1 = 8 − 3 = 5 Let the nth term of given A.P. be 78. Now as we know, an = a+(n−1)d Therefore, 78 = 3+(n −1)5 75 = (n−1)5 (n−1) = 15 n = 16 Hence, 16th term of this A.P. is 78.